# Evan Munro    Xu Kuang    Stefan Wager

### v3

@@ 5.2 Heterogeneous Treatment Effects in the Model
Optimal Rule Direct Rule Randomized Rule
@@
Table 2: Value of estimated equilibrium-stable targeting rule. Reported values are average and standard deviation across 50 samples of 5,000 households drawn from the model.
We then evaluate the optimal equilibrium stable rule by estimating the rule using the procedure in Section 4.2 on a training sample of 5,000 households, and then evaluating the expected value of that rule numerically on a separated simulated dataset of 5,000 households. The results are reported in Table 2. We compare the outcomes under this rule to two baselines. The first baseline is the randomized experiment, where 55% of eligible households are treated. The second is an equilibrium-stable version of a treatment rule that targets based on $\hat{\tau}_{\text{CADE}}(x)$ only. This rule is estimated on the training sample as $\tilde{\pi}(x)=\hat{\alpha}\cdot\mathbbm{1}(\hat{\tau}_{\text{CADE}}(x)>0)$. $\hat{\alpha}$ is chosen so that, in the training data, the total net demand under $\tilde{\pi}$ is equal to the total net demand in the randomized experiment. This rule avoids treating some of the few households for which the cash transfer is not beneficial, which improves outcomes compared to a purely randomized rule. However, for equilibrium stability, $\alpha<1$, and the rule does not distinguish between households based on their demand for the perishable protein source. The optimal rule has higher outcomes since it treats more households with a large positive direct effect on outcomes and avoids treating households for which the treatment has a small impact on outcomes but a large impact on demand.
@@ A.1 Estimation
The model described in the main text captures the impact of a conditional cash transfer on children’s health outcomes and the market for a perishable protein source in a remote village in the Philippines. It is designed based on the results in Filmer et al. (2023). The household and individual-level error terms in the demand equation have standard deviation of 1/3rd. The household level error terms in the outcome equation have standard deviation of 1/3rd and the individual-level error terms have standard deviation of 1.
@@
The model is just-identified by using these moments. Each parameter has an estimator, which is a simple function of these moments:
@@
Last, in drawing data from the model, $C_{i}$ and $h(i)$ are determined by random draws from the empirical distribution of households with children between 0 and 14 (who are in eligible for the transfer if they meet a proxy means test) from the baseline survey in Filmer et al. (2023). $E_{i}$ is drawn randomly to match the average percentage of eligible households in remote villages.
@@ B.1 Proof of Theorem 1
$P_{n}(\bm{W})-p^{*}_{\pi}=-\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi})+o_{p}(n^{-1/2}).$ (40)
and furthermore $\sqrt{n}\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)\Rightarrow\mathcal{N}\left(0,\,\operatorname{Var}\left[Z_{i}(W_{i},p^{*}_{\pi})\right]\right)$.
@@
Because $\mathcal{S}$ is compact, Assumptions 2 and 5 together imply that not only is there is a unique population equilibrium price $p^{*}_{\pi}$ with $z(p^{*}_{\pi})$, but there is also a critical radius $\varepsilon^{*}$ such that
@@
To verify this result, we note that, by twice differentiability of $z$, there exists a $\delta^{*}>0$ such that
$\left\lVert z(p)\right\rVert_{2}\geq\frac{\lambda_{\min}\left(\xi_{z}\right)}{2}\left\lVert p-p^{*}_{\pi}\right\rVert_{2}\text{ for all }\left\lVert p-p^{*}_{\pi}\right\rVert_{2}\leq\delta^{*}.$
@@
$\varepsilon^{*}=\min\left\{\delta^{*},\,\frac{2}{\lambda_{\min}\left(\xi_{z}\right)}\inf\left\{\left\lVert z(p)\right\rVert_{2}:\left\lVert p-p^{*}_{\pi}\right\rVert_{2}>\delta,\,p\in\mathcal{S}\right\}\right\},$
where the latter term must because $\mathcal{S}$ is compact, $z(p)$ is continuous over $\mathcal{S}$ and $p^{*}_{\pi}$ is the unique equilibrium point over $\mathcal{S}$.
@@
$\mathbb{P}\left[\left\lVert z(P_{n}(\bm{W}))\right\rVert>\frac{2+t}{\sqrt{n}}\right]\leq Ct^{2J}e^{-2t^{2}}+e^{-cn}.$
Together with (41), we then conclude that, for $n$ large enough, we have
$\mathbb{P}\left[\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert>\frac{1}{\sqrt{n}}\frac{4+2t}{\lambda_{\min}\left(\xi_{z}\right)}\right]\leq Ct^{2J}e^{-2t^{2}}+e^{-cn}$ (43)
for all $t\leq\sqrt{n}\lambda_{\min}\left(\xi_{z}\right)\varepsilon^{*}/2-2$, which in particular implies that $P_{n}(\bm{W})=p^{*}_{\pi}+\mathcal{O}_{p}(1/\sqrt{n})$.
@@
Finally recall that, by Assumption 3, $\operatorname{Var}\left[Z_{i}(W_{i},\,p^{*}_{\pi})\right]\succ 0$ and so weak continuity implies that $\lim_{n\rightarrow\infty}\text{Corr}\left[Z_{i}(W_{i},\,p^{*}_{\pi}),\,Z_{i}(W_{i},\,p^{*}_{\pi}+U_{i})\right]=1$. The claimed central limit theorem then follows by the standard Lindeberg argument thanks to boundedness of $Z_{i}(w,\,p)$. ∎
@@ B.2 Proof of Theorems 2, 3 and 4
$\tau_{\text{ADE}}$ $=\mathbb{E}_{\bm{\pi}}\left[\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,P_{n}(w_{i}=1;\,\bm{W}_{-i})+U_{i})-Y_{i}(0,P_{n}(w_{i}=0;\,\bm{W}_{-i})+U_{i})\right)\right]$
$=\mathbb{E}_{\bm{\pi}}\left[\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,P_{n}(\bm{W})+U_{i})-Y_{i}(0,\,P_{n}(\bm{W})+U_{i})\right)+R_{n}(1)-R_{n}(0)\right],$
@@
$\tau_{\text{ADE}}=\mathbb{E}_{\bm{\pi}}\Bigg{[}\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,p^{*}_{\pi}+U_{i})-Y_{i}(0,\,p^{*}_{\pi}+U_{i})\right)$
@@
where we note that the $o_{p}(1/\sqrt{n})$ residual term can be moved outside the $\mathbb{E}_{\pi}\left[.\right]$ term because it is almost surely bounded. Finally, plugging in the asymptotic expansion (40) for $P_{n}(\bm{W})-p^{*}_{\pi}$ into this expression gives
$\tau_{\text{ADE}}=\mathbb{E}_{\bm{\pi}}\Bigg{[}\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,p^{*}_{\pi}+U_{i})-Y_{i}(0,\,p^{*}_{\pi}+U_{i})\right)$
@@
We can now invoke Lemma 17 as well as approximate monotonicity as in Assumption 3 to check that, for any $\varepsilon>0$,
$\left\lvert R_{n}(w)\right\rvert$ $\leq C\,\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{J}\Big{(}Z_{ij}(w,\,P_{n}(\bm{W})-\varepsilon/\sqrt{n}e_{j}+U_{i})$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad-Z_{ij}(w,\,P_{n}(\bm{W})+\varepsilon/\sqrt{n}e_{j}+U_{i})\Big{)}+o_{p}(1/\sqrt{n})$
$\leq-\frac{2C\varepsilon}{\sqrt{n}}\sum_{j=1}^{J}\frac{\partial}{\partial p_{j}}z_{j}(w,\ p^{*}_{\pi})+o_{p}(1/\sqrt{n}),$
where the second line followed from asymptotic equicontinuity as used in the proof of Lemma 15. Because this bound holds for any $\varepsilon>0$, we in fact have $R_{n}(w)=o_{p}(1/\sqrt{n})$. ∎
@@
$\sup_{i=1,\ldots,\ n}\left\lVert P_{n}(w_{i}=1;\bm{W}_{-i})-P_{n}(w_{i}=0;\bm{W}_{-i})\right\rVert_{2}=o_{p}(1/\sqrt{n}).$ (51)
@@
By Assumption 1 and a union bound, we know that with probability at least $1-2ne^{-cn}$,
@@
We can thus use asymptotic equicontinuity as in the proof of Lemma 15 to argue that
$\sup_{i,w}\left\lvert P_{n}(w_{i}=w,\,\bm{W}_{-i})-p^{*}_{\pi}+\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right\rvert=o_{p}(1/\sqrt{n}),$
which implies our desired result as the asymptotic linear representation is the same regardless of $i$ and $w$. ∎
@@
$\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},P_{n}(\bm{W})+U_{i})=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+U_{i})-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+B_{n},$ (55)
@@
$\mathbb{E}_{\pi}\left[\sum\limits_{k=1}^{n}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right]=\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}\left(Z_{i}(1,p^{*}_{\pi}+U_{i})-Z_{i}(0,p^{*}_{\pi}+U_{i})\right),$
@@
$B_{n}$ $=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},P_{n}(\bm{W})+U_{i})-\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+U_{i})-\left(y_{n}\left(P_{n}(\bm{W})+U_{i})\right)-y_{n}\left(p^{*}_{\pi}\right)\right)$
$\quad\quad+y_{n}\left(P_{n}(\bm{W})+U_{i})\right)-y_{n}\left(p^{*}_{\pi}\right)-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i}).$
Furthermore, (43) in the proof of Lemma 14 gives us a tail bound for $\sqrt{n}\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert_{2}$ with an exponential tail decay. In particular, $\mathbb{P}\left[\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert_{2}>n^{-1/3}\right]$ decays exponentially fast in $n$ (because $n^{-1/3}=t_{n}/\sqrt{n}$ for $t_{n}=n^{1/6}$) so we can use the $L_{2}$-bound on the residual in Lemma 20 to verify that the second moment of the asymptotic equicontinuity residual on the first line of our decomposition decays as $o(1/n)$. Meanwhile,
$y_{n}\left(P_{n}(\bm{W})+U_{i})\right)-y_{n}\left(p^{*}_{\pi}\right)-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})$
$\quad\quad\quad\quad=y\left(P_{n}(\bm{W})+U_{i})\right)-y\left(p^{*}_{\pi}\right)-\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+o(1/\sqrt{n}),$
@@
$\left(P_{n}(\bm{W})+U_{i})-p^{*}_{\pi}\right)y^{\prime}\left(p^{*}_{\pi}\right)=\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+o(1/\sqrt{n}).$

### v4

@@ 5.2 Heterogeneous Treatment Effects in the Model
Target-Optimal Target-Direct Target-Random
@@
Table 2: The first row is the average height-for-age Z-Score of children aged 0-5 in a test sample of 5,000 households, where outcomes are simulated from the model in Section A.2 under three different treatment rules, averaged across 50 simulations. For each of the 50 simulations, the treatment rules are estimated on a training sample of 5,000 samples. The second row is the standard deviation of the average outcomes across the 50 simulations. Target-Random treats a random 55% of eligible households. Target-Optimal treats households according to treatment rule that solves an empirical version of the linear program in Section 4.2, with $\hat{\tau}_{\text{CADE}}(x)$ and $\hat{\tau}^{z}_{\text{CADE}}(x)$ estimated using a causal forest. The equilibrium constraint ensures that the equilibrium price is the same as the Target-Random rule in the training sample. Target-Direct treats eligible households according to $\tilde{\pi}(x)=\hat{\alpha}\cdot\mathbbm{1}(\hat{\tau}_{\text{CADE}}(x)>0)$. $\hat{\alpha}$ is chosen so that, in the training data, the total net demand under $\tilde{\pi}$ is equal to the total net demand under Target-Random.
We then evaluate the optimal equilibrium stable rule by estimating the rule using the procedure in Section 4.2 on a training sample of 5,000 households, and then evaluating the expected value of that rule numerically on a separated simulated dataset of 5,000 households. The results are reported in Table 2. Note that our health outcome, which is average Z-Scores of children under 5 years in the simulated village, is negative since children living in remote and isolated villages in the Philippines are smaller than average. The closer Z-scores are to zero, the more a targeting rule has improved health outcomes for young children in the simulated village.
Compared to Target-Random, Target-Direct avoids treating some of the few households for which the cash transfer is not beneficial, which improves outcomes by 8.9%. However, for equilibrium stability, $\hat{\alpha}<1$, and the rule does not distinguish between households based on their demand for the perishable protein source. Target-Optimal has improved health outcomes since it treats more households with a large positive direct effect on outcomes and avoids treating households for which the treatment has a small impact on outcomes but a large impact on demand. It improves health outcomes by 17% compared to Target-Random.
@@ A.1 Estimation
The model described in the main text captures the impact of a conditional cash transfer on children’s health outcomes and the market for a perishable protein source in a remote village in the Philippines. It is designed based on the results in Filmer et al. (2023).
@@
We use these moments to estimate the model. The randomization of treatment makes estimating coefficients on treatments straightforward, because the errors in the demand equation are independent of $W_{h(i)}$. Under the assumptions in Filmer et al. (2023), the randomization of treatment also provides enough variation to estimate coefficients on prices in a linear demand and supply model. The coefficient on price in the demand model is provided directly by the elasticity of demand taken from the literature in Filmer et al. (2023). Given the demand elasticity, and under the assumption that $W_{h(i)}$ does not enter into the supply equation, then the change in prices in treated in and control villages provides an estimate for the supply elasticity under the market-clearing condition. The authors argue that treatments do not affect supply meaningfully, because most eggs sold in small villages are produced outside each village. Formally, each parameter has an estimator, which is a simple function of the moments.
@@
Last, for simulating data from the model, $C_{i}$ and $h(i)$ are determined by random draws from the empirical distribution of households with children between 0 and 14 (who are in eligible for the transfer if they meet a proxy means test) from the baseline survey in Filmer et al. (2023). $E_{i}$ is drawn randomly to match the average percentage of eligible households in remote villages. The household and individual-level error terms in the demand equation are independent and mean-zero normally distributed variables with a standard deviation of 1/3rd. Similarly, the household level error terms in the outcome equation are independent and mean-zero normally distributed variables with a standard deviation of 1/3rd and the individual-level error terms have standard deviation of 1.
@@ B.1 Proof of Theorem 1
$P_{n}(\bm{W})-p^{*}_{\pi}=-\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+o_{p}(n^{-1/2}).$ (40)
and $\mathbb{E}\left[\left|\left|P_{n}(\bm{W})-p^{*}_{\pi})+\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right|\right|_{2}^{2}\right]=o(1/n)$. Furthermore, $\sqrt{n}\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)\Rightarrow\mathcal{N}\left(0,\,\operatorname{Var}\left[Z_{i}(W_{i},p^{*}_{\pi})\right]\right)$.
@@
Because $\mathcal{S}$ is compact, Assumptions 3 and 6 together imply that not only is there is a unique population equilibrium price $p^{*}_{\pi}$ with $z(p^{*}_{\pi})$, but there is also a critical radius $\varepsilon^{*}>0$ such that
@@
To verify this result, we note that, by twice differentiability of $z$,
$z(p)-z(p^{*}_{\pi})=\xi_{z}(p-p^{*}_{\pi})+h(p-p^{*}_{\pi})||p-p^{*}_{\pi})||$
for a continuous function $h(p-p^{*}_{\pi})$ with $h(0)=0$. Let $\lambda_{\min}\left(\xi_{z}\right)$ be the minimum eigenvalue of $\xi_{z}$ and let $\delta^{*}$ be such that $h(p-p^{*}_{\pi})\leq\lambda_{\min}\left(\xi_{z}\right)/2$ for all $||p-p^{*}_{\pi}||\leq\delta^{*}$. Then, for all $||p-p^{*}_{\pi}||\leq\delta^{*}$,
$z(p)-z(p^{*}_{\pi})\geq\lambda_{\min}\left(\xi_{z}\right)||p-p^{*}_{\pi}||_{2}-\lambda_{\min}\left(\xi_{z}\right)/2||p-p^{*}_{\pi}||_{2}=\frac{\lambda_{\min}\left(\xi_{z}\right)}{2}||p-p^{*}_{\pi}||_{2}.$
@@
$\varepsilon^{*}=\min\left\{\delta^{*},\,\frac{2}{\lambda_{\min}\left(\xi_{z}\right)}\inf\left\{\left\lVert z(p)\right\rVert_{2}:\left\lVert p-p^{*}_{\pi}\right\rVert_{2}>\delta^{*},\,p\in\mathcal{S}\right\}\right\},$
where the latter term must exist because $\mathcal{S}$ is compact, $z(p)$ is continuous over $\mathcal{S}$ and $p^{*}_{\pi}$ is the unique equilibrium point over $\mathcal{S}$.
@@
$\mathbb{P}\left[\left\lVert z(P_{n}(\bm{W}))\right\rVert>\frac{2+t}{\sqrt{n}}\right]\leq Ct^{2J}e^{-2t^{2}}+e^{-c_{1}n}.$ (43)
(43) together with (41) implies that the probability that $||z(P_{n}(\bm{W}))||_{2}\geq\epsilon^{*}$ decreases in $n$ exponentially. So (41) then implies that for $n$ large enough, that $||P_{n}(\bm{W})-p^{*}_{\pi}||_{2}\leq z||P_{n}(\bm{W}))||_{2}$ with high probability, and we have the following tail bound available for $P_{n}(\bm{w})$ for large enough $n$
$\mathbb{P}\left[\left\lVert P_{n}(\bm{W})-p^{*}_{\pi}\right\rVert>\frac{1}{\sqrt{n}}\frac{4+2t}{\lambda_{\min}\left(\xi_{z}\right)}\right]\leq Ct^{2J}e^{-2t^{2}}+e^{-c_{1}n},$ (44)
which in particular implies that $P_{n}(\bm{W})=p^{*}_{\pi}+\mathcal{O}_{p}(1/\sqrt{n})$.
@@
By the tail bound on $P_{n}(\bm{W})-p^{*}_{\pi}$ in (44), the bound on the expectation of the remainder term in Lemma 22, and boundedness of outcome and net demand, then we also have the remainder term $R_{n}=P_{n}(\bm{W})-p^{*}_{\pi}-\bar{Z}_{n}(p^{*}_{\pi})$ is such that $\mathbb{E}[||R_{n}||^{2}_{2}]=o(1/n)$. The details of this argument follow the same approach as (59).
Finally recall that, by Assumption 4, $\operatorname{Var}\left[Z_{i}(W_{i},\,p^{*}_{\pi})\right]\succ 0$ and so weak continuity implies that $\lim\limits_{n\to\infty}\text{Corr}\left[Z_{i}(W_{i},\,p^{*}_{\pi}),\,Z_{i}(W_{i},\,p^{*}_{\pi}+U_{i})\right]=1$. The claimed central limit theorem then follows by the standard Lindeberg argument thanks to boundedness of $Z_{i}(w,\,p)$. ∎
@@ B.2 Proof of Theorems 2, 3 and 4
$\tau_{\text{ADE}}$ $=\mathbb{E}_{\pi}\left[\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,P_{n}(w_{i}=1;\,\bm{W}_{-i})+U_{i})-Y_{i}(0,P_{n}(w_{i}=0;\,\bm{W}_{-i})+U_{i})\right)\right]$
$=\mathbb{E}_{\pi}\left[\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,P_{n}(\bm{W})+U_{i})-Y_{i}(0,\,P_{n}(\bm{W})+U_{i})\right)+R_{n}(1)-R_{n}(0)\right],$
@@
$\tau_{\text{ADE}}=\mathbb{E}_{\pi}\Bigg{[}\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,p^{*}_{\pi}+U_{i})-Y_{i}(0,\,p^{*}_{\pi}+U_{i})\right)$
@@
Finally, plugging in the asymptotic expansion (40) for $P_{n}(\bm{W})-p^{*}_{\pi}$ into this expression gives
$\tau_{\text{ADE}}=\mathbb{E}_{\pi}\Bigg{[}\frac{1}{n}\sum_{i=1}^{n}\left(Y_{i}(1,\,p^{*}_{\pi}+U_{i})-Y_{i}(0,\,p^{*}_{\pi}+U_{i})\right)$
@@
To finish controlling the remainder terms, we work with the two terms in this bound separately. First, by Lemma 17, we have:
$\begin{split}C\mathbb{E}_{\pi}\left[\,\frac{1}{n}\sum_{i=1}^{n}\left\lVert P_{n}(w_{i}=1;\,\bm{W}_{-i})-P_{n}(w_{i}=0;\,\bm{W}_{-i})\right\rVert_{2}\right]=o_{p}(1).\end{split}$
Next, we invoke Lemma 17 as well as approximate monotonicity as in Assumption 4. Specifically, for any $\varepsilon>0$, then Lemma 17 implies that $\left\lVert P_{n}(w_{i};\,\bm{W}_{-i})-P_{n}(\bm{W})\right\rVert_{2}\leq\varepsilon/(C\sqrt{n})$ with probability tending to 1. By Assumption 4, where $C$ is the constant from the assumption, this means that
$Z_{ij}(P_{n}(\bm{W})+\varepsilon/\sqrt{n}e_{j}+U_{i})\leq Z_{ij}(w,\,P_{n}(w_{i};\,\bm{W}_{-i})+U_{i})\leq Z_{ij}(P_{n}(\bm{W})-\varepsilon/\sqrt{n}e_{j}+U_{i})$
simultaneously for all $i$, with probability tending to 1. Plugging this into the above, we can now write
$\mathbb{E}_{\pi}[\left\lvert R_{n}(w)\right\rvert]$ $\leq\mathbb{E}_{\pi}\left[C\,\frac{1}{n}\sum_{i=1}^{n}\sum_{j=1}^{J}\Big{(}Z_{ij}(w,\,P_{n}(\bm{W})-\varepsilon/\sqrt{n}e_{j}+U_{i})-Z_{ij}(w,\,P_{n}(\bm{W})+\varepsilon/\sqrt{n}e_{j}+U_{i})\Big{)}\right]$
$\qquad\qquad\qquad\qquad+o_{p}(1/\sqrt{n}),$
$\leq-\frac{2C\varepsilon}{\sqrt{n}}\sum_{j=1}^{J}\frac{\partial}{\partial p_{j}}z_{n,j}(w,\ p^{*}_{\pi})+o_{p}(1/\sqrt{n}),$
where the second line followed from asymptotic equicontinuity of $Z_{i}(w,p+U_{i})$, from Lemma 22. Because this bound holds for any $\varepsilon>0$, we in fact have $\mathbb{E}_{\pi}[|R_{n}(w)|]=o_{p}(1/\sqrt{n})$. ∎
@@
$\sup_{i=1,\ldots,\ n}\mathbb{E}_{\pi}[\left\lVert P_{n}(w_{i}=1;\bm{W}_{-i})-P_{n}(w_{i}=0;\bm{W}_{-i})\right\rVert_{2}]=o_{p}(1/\sqrt{n}).$ (52)
@@
By Assumption 2 and a union bound, we know that with probability at least $1-2ne^{-c_{1}n}$,
@@
Since we have now shown that $\sup_{i,w}||\bar{Z}_{n}(P_{n}(w_{i}=w;\,\bm{W}_{-i}))||_{2}=o(1/\sqrt{n})$ with probability $1-2ne^{-c_{1}n}$, we can now follow (43) to (44) to show the following tail bound. For $n$ large enough, there is a finite constant $C$ such that
$\mathbb{P}\left[\sup_{i,w}||P_{n}(w_{i}=w;\,\bm{W}_{-i})-p^{*}_{\pi}||>\frac{1}{\sqrt{n}}\frac{4+2t}{\lambda_{\min}\left(\xi_{z}\right)}\right]\leq Ct^{2J}e^{-2t^{2}},$
and following the same argument as (45) to (47), we can show:
$\sup_{i,w}\left|\left|P_{n}(w_{i}=w,\,\bm{W}_{-i})-p^{*}_{\pi}+\xi_{z}^{-1}\,\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right|\right|_{2}=o_{p}(1/\sqrt{n}),$ (55)
which implies our desired result as the asymptotic linear representation is the same regardless of $i$ and $w$. We also note that as in Lemma 14, the $o_{p}(n^{-1/2})$ remainder term in this expansion, $R_{n}$, also has the property $\mathbb{E}_{\pi}[R_{n}]=o_{p}(n^{-1/2})$. ∎
@@
$\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},P_{n}(\bm{W})+U_{i})=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+U_{i})-\xi_{y}^{\top}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+B_{n},$ (57)
@@
$\mathbb{E}_{\pi}\left[\sum\limits_{k=1}^{n}\frac{W_{k}-\pi_{k}}{\pi_{k}(1-\pi_{k})}\xi_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})\right]=\xi_{y}^{\top}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}\left(Z_{i}(1,p^{*}_{\pi}+U_{i})-Z_{i}(0,p^{*}_{\pi}+U_{i})\right),$
@@
$\begin{split}B_{n}&=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},P_{n}(\bm{W})+U_{i})-\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+U_{i})-\frac{1}{n}\sum\limits_{i=1}^{n}\left(y_{n}\left(P_{n}(\bm{W})\right)-y_{n}\left(p^{*}_{\pi}\right)\right)\\ &\quad\quad+\frac{1}{n}\sum\limits_{i=1}^{n}\left(y_{n}\left(P_{n}(\bm{W})\right)-y_{n}\left(p^{*}_{\pi}\right)\right)-\xi_{y}^{\top}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i}).\end{split}$ (58)
For the first line, let
$R_{n}=\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},P_{n}(\bm{W})+U_{i})-\frac{1}{n}\sum\limits_{i=1}^{n}Y_{i}(W_{i},p^{*}_{\pi}+U_{i})-\frac{1}{n}\sum\limits_{i=1}^{n}\left(y_{n}\left(P_{n}(\bm{W})\right)-y_{n}\left(p^{*}_{\pi}\right)\right).$
Let $\delta_{n}=n^{-1/3}$. By Lemma 22, $\mathbb{E}[R^{2}_{n}\mathbbm{1}(||P_{n}(\bm{W})-p^{*}_{\pi}||_{2}\leq\delta_{n})]=o(1/n)$.
$\begin{split}\mathbb{E}\left[R_{n}^{2}\right]&\leq\mathbb{E}\left[R_{n}^{2}1\left(\left\{\left\lVert P_{n}-p^{*}_{\pi}\right\rVert_{2}\leq\delta_{n}\right\}\right)\right]+\mathbb{E}\left[R_{n}^{2}1\left(\left\{\left\lVert P_{n}-p^{*}_{\pi}\right\rVert_{2}>\delta_{n}\right\}\right)\right]\\ &\leq o(1/n)+\mathbb{P}\left[\left\lVert P_{n}-p^{*}_{\pi}\right\rVert_{2}>\delta_{n}\right]M^{2},\\ &=o(1/n).\end{split}$ (59)
where the last step is by the tail bound in (44), which indicates that $\mathbb{P}\left[\left\lVert P_{n}-p^{*}_{\pi}\right\rVert_{2}>\delta_{n}\right]=o(1/n)$. And, $M$ is finite constant since $R_{n}^{2}$ is uniformly bounded. For the second line in (58),
$\frac{1}{n}\sum\limits_{i=1}^{n}\left(y_{n}\left(P_{n}(\bm{W})\right)-y_{n}\left(p^{*}_{\pi}\right)\right)-\xi_{y}^{\top}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})$
$\quad\quad\quad\quad=\frac{1}{n}\sum\limits_{i=1}^{n}\left(y\left(P_{n}(\bm{W})\right)-y\left(p^{*}_{\pi}\right)\right)-\xi^{\top}_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+o(1/\sqrt{n}),$
@@
$\frac{1}{n}\sum\limits_{i=1}^{n}\xi_{y}^{\top}\left(P_{n}(\bm{W})-p^{*}_{\pi}\right)=\xi^{\top}_{y}\xi_{z}^{-1}\frac{1}{n}\sum\limits_{i=1}^{n}Z_{i}(W_{i},p^{*}_{\pi}+U_{i})+o_{p}(1/\sqrt{n}).$